If 6 boys and 15 men together earn $7,128 a week and if each man earns 4 times as much as each boy, how much does each boy earn in a week?
Answer by mananth(16949)

let each boy earn $x-------------6
Man earns 4x--------------------15
6x+15*4x=7128
66x=7128
/66
x=7128/66
x=$108 --- boy's earning